1. Upon analyzing a sample of silver nitrate according to Volhard’s procedure, 3.000 g of this impure sample was weighed and diluted to 100.0 mL in a volumetric flask. 25.00 mL of this solution was then titrated with a 0.05111 M KSCN requiring 35.50ml to reach the light brown color of the end point. Calculate the weight % of AgNO3
1
Expert's answer
2021-03-29T06:06:53-0400

AgNO3 + KSCN ---> AgSCN + KNO3

C1= ?, V1=25.00ml

C2= 0.05111M, V2= 35.50ml

C1V1=C2V2

C1=C2V2/V1= 0.05111x35.50/25.00=0.0725M

No of moles of AgNO3= CxV= 0.0725x100/1000= 0.00725moles

Molar mass of AgNO3= 170g/mol

Mass of AgNO3= mole x molar mass

= 0.00725x170=1.2325g

Mass of sample=3.000g

%weight of AgNO3= 1.2325/3.000 x 100%=41.08%